OpenStudy (anonymous):

Find the limit. lim t->0 <(e^-9t-1)/t,t^2/(t^3-t^2),-1/(9+t)>

OpenStudy (anonymous):

let's start with the second one.. take t^2 as a common factor you get t^2/((t^2)(t-1)) that makes the function 1/(t-1). insert zero for t and you get 1/(-1) -> -1

OpenStudy (anonymous):

for the third just insert zero. you get -1/9

OpenStudy (anonymous):

i can't answer the first part.. something is missing or you forgot some brackets.. e^(what)

myininaya (myininaya):

\[\lim_{t \rightarrow 0}\frac{e^{-9t}-1}{t} ?\]

myininaya (myininaya):

i would use l'hospital there

myininaya (myininaya):

\[\lim_{t \rightarrow 0}\frac{-9e^{-9t}-0}{1}=-9e^{-9 \cdot 0}=-9(1)=-9\]